OpenStudy (anonymous):

Select the ordered pair from the choices below that is a solution to the following system of equations: 4y = x - 13 2x + 16y = 20 A (14, 1/4) B(-3, 1) C(12, -1/4) D(5, -2) How do I do this?

OpenStudy (anonymous):

Solve the two equations 4y = x - 13 2x + 16y = 20

OpenStudy (anonymous):

and get the value of x and y

OpenStudy (anonymous):

@mattfoley ??

OpenStudy (anonymous):

can u?

OpenStudy (anonymous):

Easiest way is to solve the top equation for x then plug in x in the 2nd.

OpenStudy (anonymous):

If you do not want to solve, you try each pair and see which one works.

OpenStudy (anonymous):

Yep thats another method

OpenStudy (anonymous):

The only pair that satisfy the two equations is C. A, B and D do not work. Solving it is not also difficult.

OpenStudy (anonymous):

\[ 4y = x - 13 \\ 2x + 16y = 20\\ -x + 4y = -13\\ 2x + 16 y =20\\ -2x + 8y = -26\\ 2x + 16 y =20\\ 24 y =-6\\ y=-\frac 1 4\\ 2x - 16\frac 1 4 =20\\ 2x -4=20\\ 2x =24\\ x=12 \]